Can you please help me with this question?
nine times a number minus three is equal to five times that same number added to seven. What is the value of the number?

Answer by sudhanshu_kmr(1152)


let assume number = x

9x - 3 = 5x + 7
=> 4x = 10
=> x = 10/4

x = 5/2